A hockey puck with an initial speed of 8.0 m/s coasts 45 m to a stop across the ice. If the force of friction on the puck has a magnitude of 0.12 N, what is the puck’s mass?
1
Expert's answer
2013-01-14T11:00:05-0500
The work done by a constant force of magnitude F ona point that moves a distance d in the direction of the forceis the product,
A = F*d
The law of conservationof energy:
A = T
T - kinetic energy
T = m*v^2/2
F*d = m*v^2/2
m =2*F*d/v^2 =2*0.12*45/8^2 = 0.169 kg